28. Last year, a small nation with abundant forests cut down $200 worth of trees. $100 worth of trees were then turned into $150 worth of lumber. $100 worth of that lumber was used to produce $250 worth of bookshelves. Assuming the country produces no other outputs, and there are no other inputs used in the production of trees, lumber, and bookshelves, what is this nation's GDP? In other words, what is the value of the final goods produced including trees, lumber, and bookshelves?
Answer #1

GDP would be calculated using the value-added method:

Value added by cutting down =200-100 = 100

Value added by 150 worth of lumber = 150-100 = 50

Value added by bookshelves = 250

GDP = 100+50+250 = 400